What will be the PH of saturated soln of Mg(OH) 2 if its Ksp = 5 * 10-71)10 2)11.7 3)12 4)13

Grade:12

1 Answers

Arun
6 years ago
Dear Shreya
 
 
Mg(OH)2 (s) Mg2+ (aq) + 2 OH- (aq) 
let x = mol/L of Mg(OH)2 that dissolve : we get x mol/L Mg2+ and 2x mol/L OH- 

Ksp = 5 x 10^-7 = [Mg2+][OH-]^2 = (x)(2x)^2 = 4x^3 

x = 0.005 M 
2x = [OH-]= 0.010M 
pOH = 2 
pH = 14 - 2 =12
 
 
Regards
Arun (askIITians forum expert)